undeadknight26 (undeadknight26):

Given f(x) = (x - 7)/3, solve for f-1(4). -1 19 33 37

OpenStudy (asnaseer):

write this as: y = (x-7)/3 now rearrange this to get x in terms of y - what do you get?

undeadknight26 (undeadknight26):

x = (y - 7)/3?

OpenStudy (asnaseer):

that is not correct

OpenStudy (asnaseer):

you don't just swap the x and y

OpenStudy (asnaseer):

we start with: y = (x-7)/3 first multiply both sides by 3 - what do you get?

undeadknight26 (undeadknight26):

3y = (x-7)?

OpenStudy (asnaseer):

correct - now you can remove the braces to get: 3y = x - 7 next add 7 to both sides - what do you get?

undeadknight26 (undeadknight26):

10y =x

OpenStudy (asnaseer):

not quite you seem to have added 7y to the 3y instead of just adding 7 to it - try again

undeadknight26 (undeadknight26):

7 + 3y = x?

OpenStudy (asnaseer):

thats right - well done! :) now we swap the two sides to get: x = 7 + 3y final step is to swap the x and y variables to get the inverse function: y = 7 + 3x

OpenStudy (asnaseer):

now just plug x=4 into this

OpenStudy (xapproachesinfinity):

you just need to solve \(\large 4=\frac{x-7}{3}\)

OpenStudy (xapproachesinfinity):

the x value that you get is f inverse of (4)

OpenStudy (xapproachesinfinity):

it is the same thing it's just that getting inverse first then evaluate it at 4 takes a longer route than just solving for that equation which you get the answer quicker

OpenStudy (asnaseer):

@undeadknight26 - you need to use the method that you have been taught in class

OpenStudy (xapproachesinfinity):

:) of course use what you are taught! but once you master it be open to other ways of doing things